# Intelligence at the Edge of Chaos

### v1

@@ 3 Methodology
Figure 1: Framework for investigating the link between complexity and intelligence. We pretrain Large Language Models (LLMs) on Elementary Cellular Automata (ECAs) from different complexity classes using next-token prediction, then evaluate them on downstream reasoning and chess move prediction tasks. We use various measures to analyze the complexity of ECA rules, and quantify the relationship between complexity and downstream performance.
@@ 3.2 Training procedure for GPT-2 models
We utilized a modified GPT-2 architecture Radford et al. (2019) adapted for binary input and output data, enabling it to perform next-token prediction on sequences of binary states. Instead of using a traditional token embedding layer followed by a softmax over a vocabulary, we replaced the token embeddings with a linear projection layer that directly maps binary vectors into the model’s embedding space. The GPT-2 model processes these embeddings to capture temporal dependencies and patterns within the sequences. At the output, we apply a linear projection layer to map the model’s hidden states back to the data dimensionality, generating the prediction for the next state of binary variables at each time step. This adaptation allows the GPT-2 model to handle binary data directly and perform next-token prediction without relying on a predefined vocabulary. This also makes the model deterministic, in line with the deterministic nature of ECAs.
@@ 4 Experiments
To evaluate the emergent intelligence of models trained on cellular automata, we conducted experiments on three downstream tasks: one easy and one hard reasoning task inspired by the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019), and a challenging chess move prediction task. These tasks were designed to quantify the models’ abilities in reasoning, abstraction, and long-term prediction, thereby assessing the level of intelligence encoded during pretraining on ECA rules of varying complexities. We freeze the layers of the pretrained GPT-2 models and train only the input and output projection layers for downstream tasks to ensure that performance differences reflect the inherent capabilities of the models.
@@ 4.1 Downstream Task: Reasoning
Data Preparation and Training
For the easy reasoning task, models were trained for 1,000 epochs with a learning rate of $1\times 10^{-4}$. For the hard reasoning task, due to its increased complexity and difficulty, we train for 10,000 epochs with a learning rate of $1\times 10^{-5}$. Early stopping based on validation loss was again used to prevent overfitting.
@@ 5.1 Relationship between intelligence and complexity
We observe that models trained on certain Class III (Chaotic) rules, such as Rules 105, 146, and 150, have poorer performance on the hard reasoning and chess move prediction tasks. This behavior is explained due to chaotic systems lacking the structured patterns necessary for effective learning. In other words, they may be too random to predict, leading to weaker downstream performance. These results highlight the existence of a “sweet spot” of complexity conducive to intelligence, where the system is still predictable yet hard to predict.
@@ 5.2 Models Learn Complex Solutions For Simple Rules
Our findings reveal that models trained on rules that produce more complex dynamics tend to allocate higher attention to the last ten states, with a strong positive correlation (r=0.66) between rule complexity and average attention. This suggests that models trained on complex rules integrate information from past states to make their predictions. In contrast, models trained on simpler rules display lower attention across the last ten states. For example, Rule 168 (uniform) shows consistently low attention, indicating that previous states are not being utilized in the prediction process. Rule 179 (periodic) demonstrates a recurring pattern in its attention scores, where the model focuses on every other state. This behavior is explained by the nature of Rule 179, which produces an alternating pattern that repeats every two time steps (Figure 4). Thus, the model appears to be learning only this alternating cycle, rather than general state history. The simpler attention structure suggests that these models learn a trivial solution, which is in line with their downstream performance.
@@ 7 Discussion
8 Conclusion
We hypothesize that the emergence of intelligence is closely tied to a model’s capacity to predict and process complexity. Based on our results, we conjecture that training models on increasingly complex data may be sufficient to generate intelligent behavior. Perhaps complexity is all you need?

### v2

@@ 2.2 Large Language Models
Incorporating synthetic data into the training of large language models (LLMs) has been shown to significantly boost their performance. Synthetic data generation enables the creation of large, diverse, and high-quality datasets that may not be readily available in the real world (Anaby-Tavor et al., 2019). By augmenting training data with synthetic examples, models can learn from a broader range of linguistic patterns and rare occurrences, improving their ability to generalize to unseen scenarios (Edunov, 2018). As a result, LLMs trained with synthetic data exhibit enhanced robustness, accuracy, and adaptability across various downstream tasks (Feng et al., 2021).
@@ 3 Methodology
Figure 1: Our framework for investigating the link between complexity and intelligence. We pretrain Large Language Models (LLMs) on Elementary Cellular Automata (ECAs) from different complexity classes using next-token prediction, then evaluate them on downstream reasoning and chess move prediction tasks. We use various measures to analyze the complexity of ECA rules, and quantify the relationship between complexity and downstream performance.
@@ 3.2 Training Procedure for GPT-2 Models
We utilized a modified GPT-2 architecture (Radford et al., 2019) adapted for binary input and output data, enabling it to perform next-token prediction on sequences of binary states. Instead of using a traditional token embedding layer followed by a softmax over a vocabulary, we replaced the token embeddings with a linear projection layer that directly maps binary vectors into the model’s embedding space. The GPT-2 model processes these embeddings to capture temporal dependencies and patterns within the sequences. At the output, we apply a linear projection layer to map the model’s hidden states back to the data dimensionality, generating the prediction for the next state of binary variables at each time step. This adaptation allows the GPT-2 model to handle binary data directly and perform next-token prediction without relying on a predefined vocabulary. This also makes the model deterministic, in line with the deterministic nature of ECAs.
@@ 4 Experiments
To evaluate the emergent intelligence of models trained on cellular automata, we conducted experiments on three downstream tasks: one easy and one hard reasoning task inspired by the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019), and a challenging chess move prediction task (Ruoss et al., 2024). These tasks were designed to quantify the models’ abilities in reasoning, abstraction, and long-term prediction, thereby assessing the level of intelligence encoded during pretraining on ECA rules of varying complexities. We freeze the layers of the pretrained GPT-2 models and train only the input and output projection layers for downstream tasks to ensure that performance differences reflect the inherent capabilities of the models.
@@ 4.1 Downstream Task: Reasoning
Data Preparation and Training on the Downstream Tasks
For the easy reasoning task, models were trained for 1,000 epochs with a learning rate of $1\times 10^{-4}$. For the hard reasoning task, due to its increased complexity and difficulty, we post trained for 10,000 epochs with a learning rate of $1\times 10^{-5}$. Early stopping based on validation loss was again used to prevent overfitting.
@@ 5.1 Relationship between Intelligence and Complexity
We observe that models trained on certain Class III (Chaotic) rules, such as Rules 105, 146, and 150, have poorer performance on the hard reasoning and chess move prediction tasks. This behavior is expected due to chaotic systems lacking the structured patterns necessary for effective learning. In other words, they may be too random to predict, leading to weaker downstream performance. These results highlight the existence of a “sweet spot” of complexity conducive to intelligence, where the system is still predictable yet hard to predict.
@@ 5.2 Models Learn Complex Solutions For Simple Rules
Our findings reveal that models trained on rules that produce more complex dynamics tend to allocate higher attention to the last ten states, with a strong positive correlation (r=0.66) between rule complexity and average attention. This suggests that models trained on complex rules integrate information from past states to make their predictions. In contrast, models trained on simpler rules display lower attention across the last ten states. For example, Rule 168 (uniform) shows consistently low attention, indicating that previous states are not being utilized in the prediction process. Rule 179 (periodic) demonstrates a recurring pattern in its attention scores, where the model focuses on every other state. This behavior is expected by the nature of Rule 179, which produces an alternating pattern that repeats every two time steps (Figure 4). Thus, the model appears to be learning only this alternating cycle, rather than general state history. The simpler attention structure suggests that these models learn a trivial solution, which is in line with their downstream performance.
@@ Appendix B Long vs Short Term Prediction with other complexity measures
Figure 8 presents an efficiency comparison between models trained on short-term (1-step) and long-term (5-step) prediction tasks akin to Figure 5, but with the remaining complexity measures. We observe the same trend as with the Lempel-Ziv complexity. Models tend to perform better when trained on shorter prediction horizons, challenging the assumption that longer prediction horizons lead to better understanding of underlying structures.